# PushPull-Net: Inhibition-driven ResNet robust to image corruptions

### v1

@@ 1 Introduction
Data augmentation has been shown to considerably improve model robustness [11, 10, 41, 36]. However, mere data augmentation cannot address this problem sufficiently [41]. Improvement in the architectural design of ConvNets along with data augmentation strategies have improved model robustness [19]. We contribute to enhancing ConvNet architectural design for improved model robustness. While [19] deals with inhibiting the most significant activations, the proposed push-pull inhibition takes a contrasting approach and reduces signal responses in regions lacking preferred stimuli.
@@ 4.1 Evaluation metrics
The metrics used to evaluate model robustness are task agnostic, wherein the model performance is expressed in terms of the error made by the model. We denote by $E^{f}_{c,s}$ the error made by a trained classifier $f$, on a corruption type $c$, with a severity level $s$. Then the corruption error is defined by the following ratio as suggested in [11]:
${CE}^{f}_{c}=\biggl{(}{\sum_{s=1}^{5}{E^{f}_{c,s}}}\biggr{)}\bigg{/}\biggl{(}{\sum_{s=1}^{5}{E^{\text{ baseline}}_{c,s}}}\biggr{)}$ (9)
The errors made by the model $f$ are normalized to the errors made by a baseline model for a given corruption type $c$. Also, the mean corruption error $mCE$ of a model $f$ is the mean of the 15 corruption errors, i.e, ${CE}^{f}_{\text{Gaussian Noise}}$, ${CE}^{f}_{\text{Shot Noise}}$, …, and ${CE}^{f}_{\text{JPEG}}$.
@@ 4.2 Experimental setup
We investigate the model robustness by replacing the first layer of ResNet with the proposed PushPull-Conv. As the original ResNet architecture for ImageNet had a conv1 layer with 64 filters of size $7\times 7$ and a stride of 2, we retain the same configuration for the PushPull-Conv and explore various settings for the hyperparameters - average kernel size and the inhibition strength. As the $mCE$ scores (refer to Eq. 9) are always computed with respect to a baseline, we hereby always refer to the vanilla ResNet as the baseline. With this setting, we evaluate the $mCE$ scores of the PushPull variants keeping the corresponding vanilla ResNet as the baseline. For the experiments on the Cifar10 dataset, we modify the conv1 layer to have 64 filters with size $3\times 3$ and with a stride of 1. This change in hyperparameters was made to handle the smaller image dimensions ($32\times 32$ pixels) of that dataset.
@@ 4.3 Experiments
Table 1: Classification results. D denotes the dataset, $E$ is the top-1 error rate on the clean test set, and $mCE$ is the mean corruption error. The corruption types are Noise - (Ga)ussian, (Sh)ot, and (Im)pulse, Blur - (De)foucus, (Gl)ass, (Mo)tion, and (Zo)om, Weather - (Sn)ow, (Fr)ost, and (Fo)g, and Digital - (Br)ightness, (Co)ntrast, (El)astic, (Pi)xelate, and (Jp)eg. The push-pull variants are shown underneath the respective baselines. The baselines are marked with
@@
. Scores showing improved robustness are highlighted in green, including those below 1 before rounding to 2 decimal places. (Absolute scores are in Table 6 of the Supplementary Material)
D Model $E$ $mCE$ Noise Blur Weather Digital
@@ 4.4 Comparison with the state-of-the-art
Table 2: Comparison with the state-of-the-art. BlurConv1 [43], BlurConv2 [37], the previous attempt of PushPull [32], and another network architecture EfficientNet [33]. (Corresponding absolute scores are in Table 7 of the Supplementary Material)
D Model $E$ $mCE$ Noise Blur Weather Digital
@@
0.069 1.000 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
PushPull avg3 (ours) 0.076 0.982 0.88 0.86 0.99 0.99 0.93 1.04 0.97 1.01 0.91 1.06 1.07 1.09 0.99 0.98 0.94
PushPull avg5 (ours) 0.068 0.959 0.89 0.89 0.94 0.97 0.95 1.00 0.94 0.97 0.90 0.98 1.02 1.06 0.97 0.97 0.95
PushPull in [32] 0.069 0.973 0.94 0.92 0.93 0.97 0.94 1.01 0.97 0.99 0.91 1.02 0.99 1.07 0.99 0.97 0.97
BlurConv1 0.069 0.948 0.85 0.86 1.00 0.95 0.89 0.96 0.99 0.95 0.93 0.98 0.96 0.94 0.96 0.98 0.99
BlurConv2 0.098 1.244 1.06 1.10 1.07 1.34 1.13 1.46 1.34 1.27 1.14 1.35 1.41 1.37 1.29 1.26 1.08
@@
0.269 1.000 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
PushPull avg3 (PP3) 0.282 0.967 0.99 0.98 0.98 0.96 0.96 0.98 0.97 0.99 1.00 1.02 0.99 0.99 0.91 0.81 0.97
PushPull avg5 (PP5) 0.276 0.966 0.99 0.98 0.99 0.98 0.95 0.97 0.97 0.97 0.97 0.99 0.98 0.99 0.92 0.83 1.00
PushPull in [32] 0.269 0.996 1.00 1.00 1.02 0.99 1.00 0.99 0.99 0.99 0.98 1.00 0.98 0.99 0.99 0.99 1.03
BlurConv1 0.265 0.962 0.97 0.96 0.97 0.97 0.97 0.98 0.97 0.97 0.96 0.95 0.96 0.97 0.95 0.94 0.95
BlurConv2 0.276 0.938 0.90 0.89 0.89 0.99 0.97 0.99 0.96 1.00 0.96 0.98 1.00 1.00 0.98 0.70 0.87
BlurConv2+PP3 0.307 0.924 0.91 0.89 0.89 0.95 0.89 0.96 0.92 1.01 0.97 1.09 1.08 1.03 0.91 0.55 0.87
@@
The proposed approach complements the robustness offered by data augmentation schemes. We trained ResNet50 with data augmentation schemes including AutoAug, AugMix, PRIME, and TrivialAugment. On adding the PushPull layer to these models, the robustness improved further (see Table 2). This demonstrates our method’s compatibility and effectiveness in conjunction with current schemes. Moreover, in terms of absolute $mCE$ (see Table. 6 in the Supplementary Material), push-pull with PRIME augmentations achieves the best-ever $mCE$ score of 49.95% with ResNet50. This is a new benchmark when considering models with about 25.5M parameters.
Finally, the proposed method is evaluated with another ConvNet-based architecture, namely the most basic EfficientNet-b0 [33]. We integrated the push-pull layer in the first convolutional layer, that is, in the stem of EfficientNet. The results presented in Table 2 clearly demonstrate the robustness improvements achieved with PushPull-Conv. While further investigation into various ConvNet-based architectures is necessary, the evidence from ResNets and EfficientNet implementations shows promising outcomes. In fact, with EfficientNet we achieve robustness with negligible cost to clean error. We present a metric to quantify such a trade-off in the following section.
@@ 5.1 Clean error vs robustness
Configuration Clean error mCE
@@ 5.2 A Quantitative Metric for Assessing the Trade-off Between Robustness and Clean Accuracy
For a given model $m$, let the clean error be denoted as $E$ and the mean corruption error as $mCE$. The corresponding errors for the baseline are denoted by $E_{0}$ and $mCE_{0}$. On considering clean error, the net reduction in error rate for a model $m$ w.r.t a baseline is given by:
@@ 5.3 Robustness to high-frequency corruptions
In most experiments, a trend was observed where the PushPull models achieved higher robustness to high- and mid-frequency corruptions compared to their low-frequency counterparts. To illustrate this, we categorized the 15 corruption types from the Cifar10-C dataset into low-, mid-, or high-frequency corruptions based on the analysis of [41]. Frost, fog, brightness, and contrast are classified as low-frequency corruptions; glass-blur, motion-blur, zoom, snow, and elastic are classified into mid-frequency corruptions, and the rest as high-frequency. Table 4 shows the $mCE$ scores when the corruption types are grouped based on their frequency. The results indicate that PushPull offers more robustness to high-frequency corruptions than low frequencies. This trend is consistent with the other datasets and ResNet models.
Table 4: The averaged $mCE$ scores for Cifar10-C corruption types, grouped by corruption frequency. All PushPull models are trained with trainable inhibition $\alpha$.
@@ 6 Conclusion
PushPull-based convolutions improve the robustness of ResNet architectures to out-of-distribution image corruptions, especially those with high-frequency corruptions. This robustness is achieved without data augmentation and with a relatively small computational cost. Furthermore, this approach can be combined with data augmentation. In particular, when combined with PRIME augmentation, we achieved the best-ever reported $mCE$ of 49.95% using ResNet50 on ImageNet-C.

### v2

@@ 1 Introduction
Data augmentation has considerably improved model robustness [11, 10, 41, 36]. However, mere data augmentation cannot address this problem sufficiently [41]. Improvement in the architectural design of ConvNets along with data augmentation strategies have improved model robustness [19]. We contribute to enhancing ConvNet architectural design for improved model robustness. While [19] deals with inhibiting the most significant activations, the proposed push-pull inhibition takes a contrasting approach and reduces signal responses in regions lacking preferred stimuli.
@@ 4.1 Evaluation metrics
We use the mean corrupted error as a metric to measure robustness [11]. Let the error made by a classifier $f$, on a corruption type $c$, with a severity level $s$, be denoted by $E^{f}_{c,s}$. Then the (absolute) mean corruption error for a classifier $f$ is defined as:
$mCE=\frac{1}{|C|}\sum_{c\in C}\biggl{(}{CE}^{f}_{c}\biggr{)}=\frac{1}{|C|}\sum_{c\in C}\biggl{(}{\frac{1}{S}\sum_{s=1}^{S}{E^{f}_{c,s}}}\biggr{)}$ (9)
[…]
$mrCE=\frac{1}{|C|}\sum_{c\in C}\biggl{(}{CE}^{f}_{c}/{CE}^{b}_{c}\biggr{)}$ (10)
@@ 4.2 Experimental setup
We investigate the model robustness by replacing the first layer of ResNet with the proposed PushPull-Conv. As the original ResNet architecture for ImageNet had a conv1 layer with 64 filters of size $7\times 7$ and a stride of 2, we retain the same configuration for the PushPull-Conv and explore various settings for the hyperparameters - average kernel size and the inhibition strength. As the $mrCE$ scores (refer to Eq. 10) are always computed with respect to a baseline, we hereby always refer to the vanilla ResNet as the baseline. With this setting, we evaluate the $mrCE$ scores of the PushPull variants keeping the corresponding vanilla ResNet as the baseline. For the experiments on the Cifar10 dataset, we modify the conv1 layer to have 64 filters with size $3\times 3$ and with a stride of 1. This change in hyperparameters was made to handle the smaller image dimensions ($32\times 32$ pixels) of that dataset.
@@ 4.3 Experiments
Table 1: Classification results. D denotes the dataset, $E$ is the top-1 error rate on the clean test set. The corruption types are Noise - (Ga)ussian, (Sh)ot, and (Im)pulse, Blur - (De)foucus, (Gl)ass, (Mo)tion, and (Zo)om, Weather - (Sn)ow, (Fr)ost, and (Fo)g, and Digital - (Br)ightness, (Co)ntrast, (El)astic, (Pi)xelate, and (Jp)eg. The push-pull variants are shown underneath the respective baselines. The baselines are marked with
@@
. Scores showing improved robustness are highlighted in green, including those below 1 before rounding to 2 decimal places. The corresponding absolute corruption errors are in Table 6 of the Supplementary Material.
D Model $E$ $mCE$ $mrCE$ Noise Blur Weather Digital
@@ 4.4 Comparison with the state-of-the-art
Table 2: Comparison with the state-of-the-art. BlurConv1 [43], BlurConv2 [37], the previous attempt of PushPull [32], and with EfficientNet [33]. The corresponding absolute corruption errors are in Table 7 of the Supplementary Material.
D Model $E$ $mCE$ $mrCE$ Noise Blur Weather Digital
@@
.069 .2809 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
PushPull avg3 (ours) .076 .2681 .982 0.88 0.86 0.99 0.99 0.93 1.04 0.97 1.01 0.91 1.06 1.07 1.09 0.99 0.98 0.94
PushPull avg5 (ours) .068 .2808 .959 0.89 0.89 0.94 0.97 0.95 1.00 0.94 0.97 0.90 0.98 1.02 1.06 0.97 0.97 0.95
PushPull in [32] .069 .2984 .973 0.94 0.92 0.93 0.97 0.94 1.01 0.97 0.99 0.91 1.02 0.99 1.07 0.99 0.97 0.97
BlurConv1 .069 .2831 .948 0.85 0.86 1.00 0.95 0.89 0.96 0.99 0.95 0.93 0.98 0.96 0.94 0.96 0.98 0.99
BlurConv2 .098 .3648 1.24 1.06 1.10 1.07 1.34 1.13 1.46 1.34 1.27 1.14 1.35 1.41 1.37 1.29 1.26 1.08
@@
.269 .6670 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
PushPull avg3 (PP3) .282 .6450 .967 0.99 0.98 0.98 0.96 0.96 0.98 0.97 0.99 1.00 1.02 0.99 0.99 0.91 0.81 0.97
PushPull avg5 (PP5) .276 .6450 .966 0.99 0.98 0.99 0.98 0.95 0.97 0.97 0.97 0.97 0.99 0.98 0.99 0.92 0.83 1.00
PushPull in [32] .269 .6650 .996 1.00 1.00 1.02 0.99 1.00 0.99 0.99 0.99 0.98 1.00 0.98 0.99 0.99 0.99 1.03
BlurConv1 .265 .6260 .962 0.97 0.96 0.97 0.97 0.97 0.98 0.97 0.97 0.96 0.95 0.96 0.97 0.95 0.94 0.95
BlurConv2 .276 .6210 .938 0.90 0.89 0.89 0.99 0.97 0.99 0.96 1.00 0.96 0.98 1.00 1.00 0.98 0.70 0.87
BlurConv2+PP3 .307 .6166 .924 0.91 0.89 0.89 0.95 0.89 0.96 0.92 1.01 0.97 1.09 1.08 1.03 0.91 0.55 0.87
@@
The proposed approach complements the robustness offered by data augmentation schemes. We trained ResNet50 with data augmentation schemes including AutoAug, AugMix, PRIME, and TrivialAugment. On adding the PushPull layer to these models, the robustness improved further (see Table 2). This demonstrates our method’s compatibility and effectiveness in conjunction with current schemes. Moreover, in terms of absolute $mCE$, push-pull with PRIME augmentations achieves the best-ever $mCE$ score of 49.95% with ResNet50. This is a new benchmark when considering models with about 25.5M parameters.
Finally, the proposed method was evaluated with the EfficientNet-b0 [33] architecture. We integrated the push-pull layer in the first convolutional layer, that is, in the stem of EfficientNet. The results presented in Table 2 clearly demonstrate the robustness improvements achieved with PushPull-Conv. While further investigation into various ConvNet-based architectures is necessary, the evidence from ResNets and EfficientNet implementations shows promising outcomes. In fact, with EfficientNet we achieve robustness with negligible cost to clean error. We present a metric to quantify such a trade-off in the following section.
@@ 5.1 Clean error vs robustness
Configuration Clean error $\mathbf{mrCE}$
@@ 5.2 A Quantitative Metric for Assessing the Trade-off Between Robustness and Clean Accuracy
For a given model $m$, let the clean error be denoted by $E$ and the mean corruption error as $mCE$. The corresponding errors for the baseline are denoted by $E_{0}$ and $mCE_{0}$. On considering clean error, the net reduction in error rate for a model $m$ w.r.t a baseline is given by:
@@ 5.3 Robustness to high-frequency corruptions
In most experiments, a trend was observed where the PushPull models achieved higher robustness to high- and mid-frequency corruptions compared to their low-frequency counterparts. To illustrate this, we categorized the 15 corruption types from the Cifar10-C dataset into low-, mid-, or high-frequency corruptions based on the analysis of [41]. Frost, fog, brightness, and contrast are classified as low-frequency corruptions; glass-blur, motion-blur, zoom, snow, and elastic are classified into mid-frequency corruptions, and the rest as high-frequency. Table 4 shows the $mrCE$ scores when the corruption types are grouped based on their frequency. The results indicate that PushPull offers more robustness to high-frequency corruptions than low frequencies. This trend is consistent with the other datasets and ResNet models.
Table 4: The averaged $mrCE$ scores for Cifar10-C corruption types, grouped by corruption frequency. All PushPull models are trained with trainable inhibition $\alpha$.
@@ 6 Conclusion
PushPull-based convolutions improve the robustness of ResNet architectures to out-of-distribution image corruptions, especially those with high-frequency corruptions. This robustness is achieved without data augmentation and with a relatively small computational cost. Furthermore, this approach can be combined with data augmentation. In particular, when combined with PRIME augmentation, it achieves the best-ever reported $mCE$ of 49.95% using ResNet50 on ImageNet-C.